OpenStudy (anonymous):

can u please help me? What is the measure of one angle of a regular polygon with nine sides?

OpenStudy (anonymous):

firstly You must not forget the formula that is : (n-1)180 = sum of all the interior angles of the regular polygon, where n = no. of sides This formula only works when the polygon is regular. now n = 9 (9-1) 180 8*180 1440 degree /9 = one interior angle's measure 160 degree

OpenStudy (anonymous):

360/9 = exterior angle

OpenStudy (anonymous):

i thought the formula is (n-2)180

OpenStudy (anonymous):

Just subtract the answer to what I put from 180.

OpenStudy (anonymous):

No need to mess about with the formula (which, u r right is n-2).

OpenStudy (anonymous):

360/9 = 40 --->180-40= 140.

OpenStudy (anonymous):

Yes?No?

OpenStudy (anonymous):

yes